# Long Video Diffusion Generation with Segmented Cross-Attention and Content-Rich Video Data Curation

### v1

@@ 4.3 Implementation
Our work is built upon Allegro[48], an open-source video diffusion model with 2.8B parameters. Allegro generates high-quality videos up to 88 frames and 720p resolution from simple text input. Text inputs are handled by T5 [21] text encoder. A video caption is decoupled into five progressive sub-captions and a hidden state is separated into five segments, which correspond to the notations $N$ above. For post-processing, we adopt EMA-VFI[45] as the frame interpolation model, to further normalize the video speed and extend video length. During inference, for a single prompt from user input, we leverage GPT-4o as the refiner to generate five progressive sub-captions.
The training of Presto can be separated into two stages: Text-to-Video Pre-training, and Text-to-Video Fine-tuning. The pre-training stage is built upon the Allegro model with 88 frames and $720\times 1280$ resolution. The pre-training dataset contains 261k instances and we sample frames from videos at 6 FPS during this stage. Presto is trained for 1500 steps on 64 Nvidia H100 GPUs with a batch size of 256 and constant learning rate of 1e-4, processing a total of 384k videos. For fine-tuning, we pick the most content-diverse 47k instances from the pre-training dataset and fine-tune for another 500 steps with a batch size of 256 learning rate of 1e-4, processing a total of 128k videos.
@@
Table 1: Quantitative results of dimension performance on VBench. A higher score indicates better performance in a particular dimension. We focus on the semantic dimension suite to demonstrate our Presto is capable of generating content-rich videos with consistency.
@@ 5 Experiment
In this section, We demonstrate the long video generation capability of Presto via both quantitative and qualitative evaluation in Sec. 5.2 and Sec. 5.3 respectively. Moreover, we provide an ablation study in Sec. 5.4 on key components of our model, including training data, progressive sub-caption strategy, and Segmented Cross-Attention. We exhibit more results in Sec. C and Sec. B of Appendix, and discuss the limitations and failure cases in Sec. D.
@@ 5.1 Baseline Models
To evaluate the effectiveness of our Presto on content diversity and long-range coherence, we compare it with the state-of-the-art text-to-video models. We select the best open-source model, Allegro [48], and commercial system, Runaway Gen-3 [26], as our baseline models. We also compare with the recent MT2V method, TALC [4]. To highlight the importance of scenario coherence, we add a naive approach of “Merge Videos” in qualitative evaluation, by utilizing multiple texts to generate multiple short clips.
@@ 5.2 Quantitative Evaluation
Setup We use VBench for our automatic quantitative evaluation. VBench offers 946 official prompts to validate different aspects of generated videos and is a widely adopted benchmark in video generation methods. We directly report the results for models that exist in the VBench Leaderboard. We report several specific dimensions as well as the holistic dimensions in VBench, as shown in Tab. 1, which demonstrate the exceptional capability of generating long videos with rich content and long-range coherence while adhering to the input text.
Presto outperforms all state-of-the-art video generation models on Semantic Score. Specifically, Presto notably surpasses Allegro with $+5.5\%$, the commercial Gen-3 with $+3.3\%$, and the previous MT2V method TALC with $+34.1\%$. We attribute the performance improvement to the leverage of the progressive sub-captions generation strategy, which decouples the text input and improves the text information. Besides, we achieve a full mark in Dynamic Degree metrics, reflecting the superior ability to capture dynamics and preserve camera control. Compared with TALC which achieves a relatively high score of $98.6\%$ in Dynamic Degree, we achieve significantly better performance on all metrics, highlighting the long-range coherence aided by the meticulous data curation and Segmented Cross-Attention. For the degradation in quality scores, we hypothesize that it arises from the increased difficulty in maintaining consistency when the dynamics are complex and varied.
Figure 3: Qualitative comparison with the baselines in our user study. Our Presto can capture intricate text details and generate long videos with long-range coherence and rich content. For the first case, ours is the only method that captures the text details of “People hurry along the sidewalk”, while other methods fail to generate walking people. For the second case, our generated videos are of the largest camera motion and the best scenario coherence.
@@ 5.3 Qualitative Evaluation
Setup Evaluating the quality of generated videos is a highly subjective task, as automatic benchmarks often dis-align with human judgment. User study is a prevalent paradigm for qualitative assessment in previous work [43, 48, 47]. In this study, we collect 62 diverse text prompts encompassing a wide range of aspects, including humans, animals, landscapes, and so on. Human annotators are tasked with blindly comparing pairs of videos and making a preference judgment between two cases. To assess the enhanced content diversity achieved by our model, we evaluate three key dimensions: Scenario Diversity, Scenario Coherence, and Text-Video Adherence. We allow the tie situation when the differences are indistinguishable. We recruited 12 annotators, with each instance reviewed by three individuals, resulting in a total of 2,232 ratings. We report the win rate (%) with each method, and calculate an overall score by considering all three dimensions. as shown in Tab. 2.
Our model surpasses all baselines on the Overall Score, indicating that our Presto can generate videos with improved scenario diversity and coherence with text-following capabilities, even better than the commercial model Gen-3. Specifically, Presto excels Gen-3 in scenario diversity, slightly outperforms in text-video adherence, and closely matches in scenario coherence. For the SOTA open-source model, Allegro, Presto wins in all dimensions, especially on the scenario diversity. For TALC, Presto significantly outperforms in all metrics, further demonstrating the importance of curated data. We also consider the naive approach of utilizing multiple texts by “Merging Videos”, to serve as a reference for our metrics. Our method reaches similar results as Merging Videos on scenario diversity but significantly outperforms it in scenario coherence.
We further exhibit the real cases in our user study, as shown in Fig. 3. For the first case, our Presto is the only one that captures intricate text details of “People hurry along the sidewalk.” while all other methods fail to generate walking people. For the second case, the generated video from our Presto achieves the largest scenario motion while with the best long-range coherence. Considering Merge Video, which may generate highly diverse content, while failing to maintain consistency as five shots are totally different and unrelated.
@@ 5.4 Ablation Study
   Method Overall Score Dynamic Degree
   O(verlap) SCA 74.7 100.0
@@
   S(equential) SCA 73.7 $\downarrow$ 100.0 --
   I(solated) SCA 73.1 $\downarrow$ 100.0 --
@@
   w/o Meticulous Filtering 72.0 $\downarrow$ 97.2$\downarrow$
   Single Long Condition 71.8 $\downarrow$ 100.0 --
Table 3: Ablation results of the model design for different Segmented Cross-Attention Strategies, and LongTake-HD dataset curation. We report the performance of models with 360p resolution and 40 frames on VBench.
We now ablate on the key proposed components, including both Presto model design and LongTake-HD dataset curation. We use VBench to evaluate generated videos on two dimensions: 1) Overall Score; and 2) Dynamic Degree. As video generation methods consume huge computational resources in model training, we standardized our video generation model training to 360p resolution with 40 frames in this ablation study. Tab. 3 presents the ablation results on the automatic evaluation benchmark.
Segmented Cross-Attention (SCA) Strategy. We ablate the three strategies of SCA including Overlap Segmented Cross-Attention (OSCA), Sequential Segmented Cross-Attention (SSCA), and Isolated Segmented Cross-Attention (ISCA), as we mentioned in Sec. 4 and Fig. 2(b). All three strategies achieve 100% in Dynamic Degree score, indicating the general concept of SCA plays a significant role in capturing dynamics. OSCA achieves an excellent result of 74.7% Overall Score, which is the best version among all strategies. This design not only maintains content richness but also facilitates transitions between segments by introducing proximal overlapping. SSCA achieves a 73.7% Overall Score, slightly behind the OSCA. We attribute this degradation to the information from previous frames disrupting the interaction of later segmented latent features. ISCA achieves a 73.1% Overall Score, performing poorly in the experiments, for it prohibits the information exchange in adjacent segments. This ablation further underscores the significance of the overlapping technique in OSCA, facilitating smoother transitions between segments.
w/o Meticulous Filtering. We study the effect of our meticulously curated dataset LongTake-HD, with the same strategy of OSCA. We randomly select the same amount of pre-training data from the unfiltered video dataset, without the filtering of content diversity in Sec. 3.1 and captions in Sec. 3.2. Note that we still adopt the basic filtering for videos, including duration, speed, resolution, and low-level metrics, to ensure the basic visual quality, and ablate the effect of metrics contributing to rich content. The model with worse data achieves 72.0% on the overall score, largely lagging behind 2.7% in the same strategy with well-curated data. Moreover, the performance on Dynamic Degree will also drop by 2.8%, highlighting the effectiveness of our data curation pipeline.
Single Long Condition. To diminish the effect of the increased text length in our sub-captions, and demonstrate the effectiveness of separate modeling for each sub-caption in our Segmented Cross-Attention, we test the method of naive concatenation on text condition. Specifically, this approach directly concatenates $N$ sub-captions along the sequence length dimension, forming a single long text condition. This long condition will attend to the whole hidden states, the same as standard text-to-video generation approaches. The results show that this naive concatenation method yields the lowest Overall Score with a 2.9% drop compared to our proposed separate modeling strategy, indicating that our Segmented Cross-Attention effectively enhances the overall video quality.
@@ 6 Conclusion
We introduced Presto, a simple yet effective method for generating long-range coherent, content-rich, long videos. Our Presto achieves 78.5% and 100% on the VBench Semantic Score and Dynamic Degree, surpassing the existing SOTA video generation approaches. Presto utilizes the Segmented Cross Attention mechanism to integrate multiple texts concurrently, which can be seamlessly adopted in the existing diffusion model with DiT architecture. We also curate a high-quality video-texts dataset LongTake-HD from public datasets and sources. We leave more exploration about the attention mechanism and model structure (e.g., auto-regressive generation) for long video generation as our future work.
@@ Acknowledgement
A Details of LongTake-HD Dataset
@@
Negative Cases. We show the negative cases of keyframes and captions in Fig. 4(e) and Fig. 4(g) respectively. Blurry or unrelated keyframes are discarded, by analyzing the compressed image file size. Negative captions when LLMs refuse to respond, or with sensitive information, will be filtered out to improve the quality of captions.
@@
B More Qualitative Comparisons
@@
C Style Control and Camera Control
@@
D Limitations
Although our proposed Presto can generate long videos with long-range coherence and rich content, certain limitations remain. First, the generated videos sometimes exhibit a slight degradation in visual fidelity compared to the base model. We attribute this to the exclusive use of publicly accessible videos for training, which, while diverse and coherent, still do not match the higher quality of the private datasets leveraged by the base model. Second, in cases involving extreme scenario motion, some regions may display artifacts such as blurring or ghosting, as visualized in Fig. 6. These artifacts are likely a consequence of our model prioritizing scenario consistency and smoothness, which occasionally compromises spatial sharpness in high-motion backgrounds. Last, our model is not suitable for generating still frames.

### v2

@@ 4.3 Implementation
Our work is built upon Allegro[54], an open-source video diffusion model with 2.8B parameters. Allegro generates high-quality videos up to 88 frames and 720p resolution from simple text input. Text inputs are handled by T5 [26] text encoder. A video caption is decoupled into five progressive sub-captions and a hidden state is separated into five segments, corresponding to the notations $N$ above. For post-processing, we adopt EMA-VFI[51] as the frame interpolation model, to further normalize the video speed and extend video length. During inference, for a single prompt from user input, we leverage GPT-4o as the refiner to generate five progressive sub-captions.
The training of Presto can be separated into two stages: Text-to-Video Pre-training, and Text-to-Video Fine-tuning. The pre-training stage is built upon the Allegro model with 88 frames and $720\times 1280$ resolution. The pre-training dataset contains 261k instances, and we sample frames from videos at 6 FPS during this stage. Presto is trained for 1500 steps on 64 Nvidia H100 GPUs with a batch size of 256 and constant learning rate of 1e-4, processing a total of 384k videos. For fine-tuning, we pick the most content-diverse 47k instances from the pre-training dataset and fine-tune for another 500 steps with a batch size of 256 learning rate of 1e-4, processing a total of 128k videos.
@@
Table 2: Quantitative results of dimension performance on VBench. The bold means the best and the underline means the second. We focus on the semantic dimension suite to demonstrate our Presto is capable of generating content-rich videos with consistency.
@@ 5 Experiment
In this section, we demonstrate the long video generation capability of Presto via both quantitative and qualitative evaluation in Sec. 5.2 and Sec. 5.3, respectively. Moreover, we provide an ablation study in Sec. 5.4 on key components of our model, including training data, progressive sub-caption strategy, and Segmented Cross-Attention. We exhibit more results in Appendix D and Appendix E, and discuss the limitations and failure cases in Appendix F.
@@ 5.1 Baseline Models
To evaluate the effectiveness of our Presto on content diversity and long-range coherence, we compare it with the state-of-the-art text-to-video models. We select the best open-source model, Allegro [54], and commercial system, Runaway Gen-3 [31], as our baseline models. We also compare with the recent MT2V method, TALC [4]. To highlight the importance of scenario coherence, we add a naive approach of “Merge Videos" in qualitative evaluation, by utilizing multiple texts to generate multiple short clips.
@@ 5.2 Quantitative Evaluation
Setup We use VBench for our automatic quantitative evaluation. VBench offers 946 official prompts to validate different aspects of generated videos and is a widely adopted benchmark in video generation methods. To align with the evaluation of other models, we assess the original videos with 88 frames before interpolation for our Presto.
We directly report the results for models that exist in the VBench Leaderboard. We report several specific dimensions and the holistic dimensions in VBench, as shown in Tab. 2, demonstrating the exceptional capability of generating long videos with rich content and long-range coherence while adhering to the input text.
Presto outperforms all state-of-the-art video generation models on Semantic Score. Specifically, Presto notably surpasses Allegro with $+5.5\%$, the commercial Gen-3 with $+3.3\%$, and the previous MT2V method TALC with $+34.1\%$. We attribute the performance improvement to the progressive sub-captions generation strategy, which decouples the text input and improves the text information. Besides, we achieve a full mark in Dynamic Degree metrics, reflecting the superior ability to capture dynamics and preserve camera control. Compared with TALC, which achieves a relatively high score of $98.6\%$ in Dynamic Degree, we achieve significantly better performance on all metrics, highlighting the long-range coherence aided by the meticulous data curation and Segmented Cross-Attention. For the degradation in quality scores, we hypothesize that it arises from the increased difficulty in maintaining consistency when the dynamics are complex and varied.
We explore more details of the hypotheses made within this sub-section in the Appendix, including ‘progressive sub-captions contribute to higher semantics’ in Appendix B, and ‘complex dynamics lead to degradation in quality’ in Appendix C.
Figure 3: Qualitative comparison with the baselines in our user study. Our Presto can capture intricate text details and generate long videos with long-range coherence and rich content. For the first case, ours is the only method that captures the text details of “People hurry along the sidewalk", while other methods fail to generate walking people. For the second case, our generated videos are of the largest camera motion and the best scenario coherence.
@@ 5.3 Qualitative Evaluation
Setup Evaluating the quality of generated videos is a highly subjective task, as automatic benchmarks often dis-align with human judgment. A user study is a prevalent paradigm for qualitative assessment in previous work [49, 54, 53]. In this study, we collected 62 diverse text prompts covering a wide range of aspects, including humans, animals, landscapes, and more. Human annotators are tasked with blindly comparing pairs of videos and making a preference judgment between two cases. To assess the enhanced content diversity achieved by our model, we evaluate three key dimensions: 1) Scenario Diversity, which video is more diverse considering the changing scenario; 2) Scenario Coherence, which video is more coherent on maintaining objects and background in the changing scenario; and 3) Text-Video Adherence, which video is closer to the user prompt. We allow the tie situation when the differences are indistinguishable. We recruited 12 annotators, and each instance was reviewed by three individuals, resulting in 2,232 ratings. We report the win rate (%) with each method and calculate an overall score of three dimensions in Tab. 3.
Our model surpasses all baselines on the Overall Score, indicating that our Presto can generate videos with improved scenario diversity and coherence with text-following capabilities, even better than the commercial model Gen-3. Specifically, Presto excels Gen-3 in scenario diversity, slightly outperforms in text-video adherence, and closely matches in scenario coherence. For the SOTA open-source model, Allegro, Presto wins in all dimensions, especially on the scenario diversity. For TALC, Presto significantly outperforms in all metrics, further demonstrating the importance of curated data. We also consider the naive approach of utilizing multiple texts by “Merging Videos", to serve as a reference for our metrics. Our method reaches similar results as Merging Videos on scenario diversity but significantly outperforms it in scenario coherence.
We further exhibit the real cases in our user study, as shown in Fig. 3. For the first case, our Presto is the only one that captures intricate text details of “People hurry along the sidewalk." while all other methods fail to generate walking people. For the second case, our Presto achieves the largest scenario motion while with the best long-range coherence. Merge video may generate highly diverse content, but it fails to maintain consistency as five shots are different and unrelated.
@@ 5.4 Ablation Study
We now ablate on the key proposed components, including both model design and dataset curation. We use VBench to evaluate generated videos on three dimensions: 1) Overall Score; 2) Dynamic Degree; and 3) Overall Consistency. As video generation methods consume huge computational resources in model training, we standardized our video generation model training to 360p resolution with 40 frames in the ablation study. Tab. 4 presents the ablation results on the automatic evaluation benchmark.
Segmented Cross-Attention (SCA) Strategy. We ablate the three strategies of SCA, including Overlap Segmented Cross-Attention (OSCA), Sequential Segmented Cross-Attention (SSCA), and Isolated Segmented Cross-Attention (ISCA), as we mentioned in Sec. 4 and Fig. 2(b). All three strategies achieve 100% in the Dynamic Degree score, indicating that the general concept of SCA plays a significant role in capturing dynamics. OSCA achieves an excellent result of 74.7% Overall Score and 25.29% Overall Consistency, which is the best version among all strategies. This design not only maintains content richness but also facilitates transitions between segments by introducing proximal overlapping. SSCA is slightly behind OSCA, and we attribute this degradation to the information from previous frames disrupting the interaction of later segmented latent features. ISCA performs poorly in experiments, as the isolated design prohibits information exchange in adjacent segments. This ablation further underscores the significance of the overlapping technique in OSCA, facilitating smoother transitions between segments.
   Method Overall Dynamic Consistency
   O(verlap) SCA 74.7 100.0 25.29
@@
   S(equential) SCA 73.7 $\downarrow$ 100.0 25.06 $\downarrow$
   I(solated) SCA 73.1 $\downarrow$ 100.0 24.88 $\downarrow$
@@
   w/o Meticulous Filtering 72.0 $\downarrow$ 97.2 $\downarrow$ 24.06 $\downarrow$
   Single Long Condition 71.8 $\downarrow$ 100.0 24.06 $\downarrow$
Table 4: Ablation results of model design and dataset curation.
w/o Meticulous Filtering. We study the effect of our meticulously curated dataset LongTake-HD, with the same strategy of OSCA. We randomly select the same amount of pre-training data from the video datasets without the filtering of content diversity in Sec. 3.1 and captions in Sec. 3.2. Note that we still adopt basic filtering for videos, including duration, speed, resolution, and low-level metrics, to ensure basic visual quality and reduce the effect of metrics contributing to rich content. The model with worse data achieves 72.0% on the overall score, largely lagging behind 2.7% in the same strategy with well-curated data. Moreover, the performance on Dynamic Degree and Overall Consistency will also drop by 2.8% and 1.23%, respectively, highlighting the effectiveness of our data curation.
Single Long Condition. To diminish the effect of the increased text length in our sub-captions and demonstrate the effectiveness of separate modeling in our Segmented Cross-Attention, we test the method of naive concatenation on text condition. Specifically, this approach directly concatenates $N$ sub-captions along the sequence length dimension, forming a single long text condition. This long condition will attend to the whole hidden states, the same as standard text-to-video generation approaches. Results show that this naive concatenation method yields the lowest Overall Score with a 2.9% drop, indicating that our Segmented Cross-Attention effectively enhances the overall video quality.
@@ 6 Conclusion
We introduced Presto, a simple yet effective method for generating long-range coherent, content-rich, long videos. Our Presto achieves 78.5% and 100% on the VBench Semantic Score and Dynamic Degree, surpassing the existing SOTA video generation approaches. Presto utilizes the Segmented Cross Attention mechanism to integrate multiple texts concurrently, which can be seamlessly adopted in the existing diffusion model with DiT architecture. We also curate a high-quality video-texts dataset LongTake-HD from public sources. We demonstrate that high quality video-text pairs are crucial for long video generation, and our curated LongTake-HD is a good candidate for future research. We leave more exploration about the attention mechanism and model structure (e.g., auto-regressive generation) for long video generation as our future work.
@@ Acknowledgement
Appendix A Details of LongTake-HD Dataset
@@
Negative Cases. We show the negative cases of keyframes and captions in Fig. 4(e) and Fig. 4(g) respectively. Blurry or unrelated keyframes are discarded, by analyzing the compressed image file size. Negative captions with sensitive information or when LLMs refuse to respond will be filtered out to improve the quality of captions.
@@
Sub-captions Similarity $\uparrow$ ROUGE-L $\uparrow$ BLEU-4 $\uparrow$
Vanilla 0.6408 0.1968 0.0376
Progressive 0.7778 0.2306 0.0578
Table 6: Text similarity of training captions and inference captions, compared between vanilla style and progressive style.
Appendix B Details of Progressive Sub-captions
Progressive sub-captions have been demonstrated to improve semantic scores in diffusion model training [38]. Intuitively, the progressive style enhances caption coherence, mitigating redundant information and phrasing. This section offers a unique perspective to further substantiate this argument: the LLM-refined progressive style outperforms the non-refined vanilla style for training sub-captions. We adopt a text-centric approach, evaluating this hypothesis by computing the text similarity between training and inference captions (note that inference captions remain constant). This experimental design comes from the intuitive notion that closer distribution between inference and training data will yield better results. We employ Cosine Similarity [29, 28], Rouge-L [17, 16], and BLEU-4 [23] metrics to assess the text similarity. As evidenced in Tab. 6, progressive-style captions exhibit improved text similarity compared to vanilla-style captions across all metrics, indicating a better semantic score in the generated videos. This observation indirectly validates our hypothesis.
We acknowledge that the most direct validation would involve training diffusion models under identical settings with both captioning styles and subsequently comparing the quality of the generated videos. However, given the substantial computational resources required for diffusion model training, we reserve this comprehensive evaluation for future work.
Appendix C Analysis of Videos with Complex Dynamics
This section mainly analyzes the reasons for the quality degradation in videos with complex dynamics. We refer to the issue of high-dynamism training videos suffering from quality degradation
This section primarily investigates the underlying causes of quality degradation observed in videos exhibiting complex dynamics. We term the phenomenon of highly dynamic training videos experiencing quality degradation as ‘dynamism loss’. Several factors contribute to this effect: 1) Individual frames within dynamically complex videos are inherently more susceptible to motion blur; and 2) Video format compression, specifically H.264 encoding employed in our experiments, induces greater quality loss in videos with higher dynamic range. This ‘dynamism loss’ happens twice when generating a content-rich video in our experiments, occurring both during the filtering and transcoding of training video data, and subsequently during the saving and encoding of generated videos. This two-fold occurrence accounts that it’s harder to maintain the same level of video quality compared with dynamic videos and normal videos, thus explaining the observed quality decline in our generated videos.
Appendix D More Qualitative Comparisons
@@
Appendix E Style Control and Camera Control
@@
Appendix F Limitations
Although our proposed Presto can generate long videos with long-range coherence and rich content, certain limitations remain. First, the generated videos sometimes slightly degrade visual fidelity compared to the base model. We attribute this to the exclusive use of publicly accessible videos for training, which, while diverse and coherent, still do not match the higher quality of the private datasets leveraged by the base model. Second, in cases involving extreme scenario motion, some regions may display artifacts such as blurring or ghosting, as visualized in Fig. 6. These artifacts are likely a consequence of our model prioritizing scenario consistency and smoothness, which occasionally compromises spatial sharpness in high-motion backgrounds. Last, our model is not suitable for generating still frames.